A boy is b years old and his father is 6times as old as him. Find the fathers age. and also the sum of thier ages in y years time.
Answer by ReadingBoosters(3246)

b = 6b
Father is 6b, without b, or boy's age, cannot determine father's exact age.
But, if boy is say 10, then father is 6(10)=60
b+y = 6b+y
In y years, they both will have increased in age by y years.
If, for example, we look in 12 years, then
b+12=6b+12 and if, for example, boy was 10, then
10+12=6(10)+12
boy is 22 and father is 60+12= 72
........................
Delighted to help.
-Reading Boosters
Wanting for others what we want for ourselves.